# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ C.1. The jumps of $\mathbf{Q}_{n}$ near $b_{j}$
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{+}(z;b_{j})\begin{bmatrix}1&0\\ \,\mathrm{e}^{\mathrm{i}\alpha_{j}\pi}W(z;b_{j})^{-2}\,\mathrm{e}^{-n\mathfrak{g}(z)}&1\end{bmatrix}.$
@@
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-n\mathfrak{g}(z)}&1\end{bmatrix},\quad z\in\Sigma_{b_{j},1}.$
@@
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-n\mathfrak{g}(z)}&1\end{bmatrix},\quad z\in\Sigma_{b_{j},3}.$
@@ C.2. The jumps of $\mathbf{A}_{n}$ near $b_{j}$
×$\times\mathbf{E}^{-1}\left(\frac{2}{\pi n}\right)^{-\frac{1}{2}{\boldsymbol{\sigma}_{3}}}\mathrm{i}^{\boldsymbol{\sigma}_{3}}\left(\varphi(z;b_{j})_{-}^{\frac{1}{4}}\right)^{\boldsymbol{\sigma}_{3}}.$
@@
$=\mathbf{A}_{n}^{-}(z;b_{j})\Psi(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi(z)^{\boldsymbol{\sigma}_{3}},$
@@ C.3. The jumps of $\mathbf{P}_{n}$ near $b_{j}$
$=\mathbf{A}_{n}^{-}(z;b_{j})\Psi_{-}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi_{-}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{Q}^{-}_{n}(z;b_{j})\rho(z;b_{j})^{{\boldsymbol{\sigma}_{3}}}\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$
$=\mathbf{P}_{n}^{-}(z;b_{j})\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi_{-}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi_{-}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{Q}^{-}_{n}(z;b_{j})\rho(z;b_{j})^{{\boldsymbol{\sigma}_{3}}}\begin{bmatrix}0&1\\ -1&0\end{bmatrix}.$
@@ C.4. Local estimates
We have the following estimates which we state in a series of lemmas. We use the notation
@@
$\|\mathbf{Q}_{n}^{-}(z;b_{j})W(z;b_{j})^{\boldsymbol{\sigma}_{3}}\|=\mathrm{O}(\log n|z-b_{j}|),\quad n^{2}(z-b_{j})\to 0.$
@@
$\operatorname{Re}\mathfrak{g}(z)=O(|z-b_{j}|^{1/2}),\quad z\to b_{j}.$
@@
To prove this lemma we use the following lemma [^5] The proof in [14] was for $\alpha_{j}=\beta_{j}=1/2$ but the same calculations hold in general. of [14, Lemma A.2] which shows that $\,\mathrm{e}^{G(z){\boldsymbol{\sigma}_{3}}}W(z;b_{j})^{\boldsymbol{\sigma}_{3}}$ is bounded in a neighborhood of $z=b_{j}$.
@@
We have for $|z-b_{j}|\leq\epsilon$ and some $d\in\mathbb{R}$
$|\varphi(z;b_{j})-d(z-b_{j})|$ $\leq|z-b_{j}|^{2},$
$|n^{2}\varphi(z;b_{j})|-|dn^{2}(z-b_{j})|$ $\leq(n^{4}|z-b_{j}|^{2})n^{-2}\leq C^{2}/n^{2}.$
@@
where $r(z)$ is analytic in a neighborhood of $[a_{j},b_{j}]$ and is positive on $[a_{j},b_{j}]$. We have
$\operatorname{Re}\mathfrak{g}(z)=\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y^{\prime})\mathop{}\!\mathrm{d}y^{\prime}.$
@@
$\log|\varphi(M(z;a_{j},b_{j})|=\int_{0}^{y}\operatorname{Im}\frac{\mathop{}\!\mathrm{d}y^{\prime}}{\sqrt{x+\mathrm{i}y^{\prime}-b_{j}}\sqrt{x+\mathrm{i}y^{\prime}-a_{j}}}.$
@@
This results in, for $h(x+\mathrm{i}y)\geq c$,
$\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y)$ $\geq c\operatorname{Im}\frac{1}{\sqrt{x+\mathrm{i}y-b_{j}}\sqrt{x+\mathrm{i}y-a_{j}}}-C^{\prime\prime}\sqrt{y},$
$\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y)$ $\geq c\log|\varphi(M(z;a_{j},b_{j}))|-\frac{2C^{\prime\prime}}{3}y^{3/2}.$
We then claim for $0<y<\gamma$ and $a_{j}\leq x\leq b_{j}$, that there exists $c^{\prime}>0$ such that
$I(x,y)=\operatorname{Im}\frac{1}{\sqrt{x+\mathrm{i}y-b_{j}}\sqrt{x+\mathrm{i}y-a_{j}}}\geq c^{\prime}.$
@@
$\log|\varphi(M(z;a_{j},b_{j}))|\geq c^{\prime}y,$
$-y^{3/2}\geq-\frac{\sqrt{y}}{c^{\prime}}\log|\varphi(M(z;a_{j},b_{j}))|.$
@@
$\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y)$ $\geq\left(c-\sqrt{y}\frac{2C^{\prime\prime}}{3c^{\prime}}\right)\log|\varphi(M(z;a_{j},b_{j}))|,$
from which the claim follows for sufficiently small $\gamma$. Similar considerations work for $\operatorname{Im}y<0$, establishing the claim for $a_{j}\leq x\leq b_{j}$. Continuity of the ratio $\operatorname{Re}\mathfrak{g}(z)/\log|\varphi(M(z;a_{j},b_{j}))|$ near $a_{j},b_{j}$ allows the desired bound to hold in neighborhoods of $a_{j},b_{j}$, completing the proof.
@@ Appendix D Hyperelliptic Riemann surfaces
Classic references for what follows are [7, 13, 3]. Much of what follows here was closely adapted from that in [15]. Using the intervals $[a_{j},b_{j}],1\leq j\leq g+1,$ we define a Riemann surface by the zero locus of
@@
$\oint_{{\mathfrak{a}}_{i}}\sum_{k=1}^{g}c_{k}\mathop{}\!\mathrm{d}\nu_{k}=\sum_{k=1}^{g}c_{k}\mathbf{A}_{ik}=\mathbf{e}_{i}^{T}\mathbf{A}c=\mathbf{e}_{i}^{T}\mathbf{e}_{j}=\delta_{ij}.$
@@
$\frac{\theta\left(\pm\mathbf{u}(z)+{\boldsymbol{\tau}}\mathbf{e}_{j}+\mathbf{v}-\mathbf{d};{\boldsymbol{\tau}}\right)}{\theta\left(\pm\mathbf{u}(z)+{\boldsymbol{\tau}}\mathbf{e}_{j}-\mathbf{d};{\boldsymbol{\tau}}\right)}=\,\mathrm{e}^{\pm v_{k}}\frac{\theta\left(\pm\mathbf{u}(z)+\mathbf{v}-\mathbf{d};{\boldsymbol{\tau}}\right)}{\theta\left(\pm\mathbf{u}(z)-\mathbf{d};{\boldsymbol{\tau}}\right)}.$
@@
where $\mathcal{C}$ is a clockwise-oriented simple contour that encircles $[a_{1},b_{g+1}]$. Because all the differentials $\mathop{}\!\mathrm{d}\omega_{j}$ are of the form $P(z)/R(z)$ where $P$ is a degree $g-1$ polynomial and $R(z)=\mathrm{O}(z^{g+1})$ as $z\to\infty$, it follows that $\oint_{\mathcal{C}}\mathop{}\!\mathrm{d}\omega_{j}=0$. Thus, ignoring any poles $\boldsymbol{\Theta}$ may have, we find that $\boldsymbol{\Theta}$ satisfies the following jump conditions:
$\boldsymbol{\Theta}^{+}(z)=\begin{cases}\boldsymbol{\Theta}^{-}(z)\begin{bmatrix}0&1\\ 1&0\end{bmatrix}&z\in(a_{j},b_{j}),\vspace{3pt}\\ \boldsymbol{\Theta}^{-}(z)\begin{bmatrix}\,\mathrm{e}^{-v_{j}}&0\\ 0&\,\mathrm{e}^{v_{j}}\end{bmatrix}&z\in(b_{j},a_{j+1}),\\ \boldsymbol{\Theta}^{-}(z)&z\in(-\infty,a_{1})\cup(b_{g+1},\infty).\end{cases}$
@@
From [18] (see also [33, Lemma 11.10]), these divisors are nonspecial so that the $\theta$ functions we consider do not vanish identically.
@@
It is important to note that (D.5) was first used in [11] and subsequently by many, see [10, 13, 8], for example.
@@ Appendix E Estimates of some exponential integrals
$E_{\rho}$ =$=\{z~{}:~{}|z+\sqrt{z+1}\sqrt{z-1}|<\rho\},\quad\rho>1,$
@@
$I_{m}:=\int_{E_{\rho}}$ $|z^{2}-1|^{-\gamma}|\varphi(z)|^{-m}\mathop{}\!\mathrm{d}A(z).$
We then set $z=1/2(w+w^{-1})$ for $1<|w|<\rho$ and change variables. If $w=\alpha+\mathrm{i}\beta$ then the change of area element is given by
$\mathop{}\!\mathrm{d}A(z)=\left[\frac{\partial x}{\partial\alpha}\frac{\partial y}{\partial\beta}-\frac{\partial x}{\partial\beta}\frac{\partial y}{\partial\alpha}\right]\mathop{}\!\mathrm{d}A(w).$
From the Cauchy–Riemann equations, we have
$\frac{\partial x}{\partial\alpha}=\frac{\partial y}{\partial\beta},\quad\frac{\partial y}{\partial\alpha}=-\frac{\partial x}{\partial\beta},$
and therefore
$\mathop{}\!\mathrm{d}A(z)=\left|\frac{\partial z}{\partial\alpha}\right|^{2}\mathop{}\!\mathrm{d}A(w).$
And so, we compute
$\frac{\partial z}{\partial\alpha}$ $=\frac{1}{2}+\frac{1}{2}\frac{\alpha^{2}+\beta^{2}-2\alpha(\alpha-\mathrm{i}\beta)}{(\alpha^{2}+\beta^{2})^{2}}=\frac{1}{2}+\frac{1}{2}\frac{\beta^{2}-\alpha^{2}+2\mathrm{i}\alpha\beta}{(\alpha^{2}+\beta^{2})^{2}}=\frac{1}{2}+\frac{1}{2}\frac{(\beta+\mathrm{i}\alpha)^{2}}{(\alpha^{2}+\beta^{2})^{2}}=\frac{1}{2}+\frac{1}{2}\frac{(\mathrm{i}\bar{w})^{2}}{|w|^{4}}$
$=\frac{1}{2}-\frac{1}{2}\frac{\bar{w}^{2}}{|w|^{4}}=\frac{1}{2}\left[1-\frac{\bar{w}}{|w|^{2}}\right]\left[1+\frac{\bar{w}}{|w|^{2}}\right].$
Then we write
$1\pm\frac{\bar{w}}{w\bar{w}}=\frac{w\bar{w}\pm\bar{w}}{w\bar{w}}=\frac{w\pm 1}{w},$
to find
$\left|\frac{\partial z}{\partial\alpha}\right|=\frac{1}{2}\frac{|w^{2}-1|}{|w|^{2}},$
or
@@
$I_{m}=\int_{1<|w|<\rho}|w|^{2\gamma-4-m}2^{\gamma-2}|w^{2}-1|^{2-2\gamma}\mathop{}\!\mathrm{d}A(w).$

### v4

@@ C.1. The jumps of $\mathbf{Q}_{n}$ near $b_{j}$
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}1&0\\ \,\mathrm{e}^{\mathrm{i}\alpha_{j}\pi}W(z;b_{j})^{-2}\,\mathrm{e}^{-{2}n\mathfrak{g}(z)}&1\end{bmatrix}$.
@@
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-{2}n\mathfrak{g}(z)}&1\end{bmatrix},\quad z\in\Sigma_{b_{j},1}.$
@@
$\mathbf{Q}_{n}^{+}(z;b_{j})=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-{2}n\mathfrak{g}(z)}&1\end{bmatrix},\quad z\in\Sigma_{b_{j},3}.$
@@ C.2. The jumps of $\mathbf{A}_{n}$ near $b_{j}$
$\times\mathbf{E}^{-1}\left(\frac{2}{\pi n}\right)^{-\frac{1}{2}{\boldsymbol{\sigma}_{3}}}\mathrm{i}^{\boldsymbol{\sigma}_{3}}\left(\varphi(z;b_{j})_{-}^{\frac{1}{4}}\right)^{\boldsymbol{\sigma}_{3}}.$
@@
$=\mathbf{A}_{n}^{-}(z;b_{j})\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}},$
@@ C.3. The jumps of $\mathbf{P}_{n}$ near $b_{j}$
$=\mathbf{A}_{n}^{-}(z;b_{j})\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{Q}^{-}_{n}(z;b_{j})\rho(z;b_{j})^{{\boldsymbol{\sigma}_{3}}}\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$
$=\mathbf{P}_{n}^{-}(z;b_{j})\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{Q}^{-}_{n}(z;b_{j})\rho(z;b_{j})^{{\boldsymbol{\sigma}_{3}}}\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$.
@@ C.4. Local estimates
We have the following estimates which we state in a series of lemmas.
@@
$\|\mathbf{Q}_{n}(z;b_{j})W(z;b_{j})^{\boldsymbol{\sigma}_{3}}\|=\mathrm{O}(\log n|z-b_{j}|),\quad n^{2}(z-b_{j})\to 0.$
@@
$\operatorname{Re}\mathfrak{g}(z)=\mathrm{O}(|z-b_{j}|^{1/2}),\quad z\to b_{j}.$
@@
To prove this lemma we use the following lemma [^6] The proof in [17] was for $\alpha_{j}=\beta_{j}=1/2$ but the same calculations hold in general. of [17, Lemma A.2] which shows that $\mathrm{e}^{G(z){\boldsymbol{\sigma}_{3}}}W(z;b_{j})^{\boldsymbol{\sigma}_{3}}$ is bounded in a neighborhood of $z=b_{j}$.
@@
We have for $|z-b_{j}|\leq\epsilon$ and some $d\in\mathbb{R}$, ${D^{\prime}}>0$
$|\varphi(z;b_{j})-d(z-b_{j})|$ $\leq{D^{\prime}}|z-b_{j}|^{2},$
$|n^{2}\varphi(z;b_{j})|-|dn^{2}(z-b_{j})|$ $\leq(n^{4}{D^{\prime}}|z-b_{j}|^{2})n^{-2}\leq{D^{\prime}}C^{2}/n^{2}.$
@@
where $r(z)$ is analytic in a neighborhood of $a_{j},b_{j}]$ and is positive on $a_{j},b_{j}]$. We have
$\operatorname{Re}\mathfrak{g}(z)={-}\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y^{\prime})\mathop{}\!\mathrm{d}y^{\prime}.$
@@
$\log|{\phi}(M(z;a_{j},b_{j}))|={-}\int_{0}^{y}\operatorname{Im}\frac{\mathop{}\!\mathrm{d}y^{\prime}}{\sqrt{x+\mathrm{i}y^{\prime}-b_{j}}\sqrt{x+\mathrm{i}y^{\prime}-a_{j}}}.$
@@
This results in, for $c^{\prime}\geq h(x+\mathrm{i}y)\geq c$,
$c^{\prime}\operatorname{Im}\frac{1}{\sqrt{x+\mathrm{i}y-b_{j}}\sqrt{x+\mathrm{i}y-a_{j}}}+C^{\prime\prime}\sqrt{y}\geq\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y){\mathop{}\!\mathrm{d}y}$ $\geq c\operatorname{Im}\frac{1}{\sqrt{x+\mathrm{i}y-b_{j}}\sqrt{x+\mathrm{i}y-a_{j}}}-C^{\prime\prime}\sqrt{y},$
$c^{\prime}\log|{\phi}(M(z;a_{j},b_{j}))|+\frac{2C^{\prime\prime}}{3}y^{3/2}\geq\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y)\mathop{}\!\mathrm{d}y$ $\geq c\log|{\phi}(M(z;a_{j},b_{j}))|-\frac{2C^{\prime\prime}}{3}y^{3/2}.$
We then claim for $0<y<\gamma$ and $a_{j}\leq x\leq b_{j}$, that there exists ${c^{\prime\prime}}>0$ such that
$I(x,y)=\operatorname{Im}\frac{1}{\sqrt{x+\mathrm{i}y-b_{j}}\sqrt{x+\mathrm{i}y-a_{j}}}\geq{c^{\prime\prime}}.$
@@
$\log|{\phi}(M(z;a_{j},b_{j}))|\geq{c^{\prime\prime}}y,$
$-y^{3/2}\geq-\frac{\sqrt{y}}{{c^{\prime\prime}}}\log|{\phi}(M(z;a_{j},b_{j}))|.$
@@
$\left(c^{\prime}+\sqrt{y}\frac{2C^{\prime\prime}}{3{c^{\prime\prime}}}\right)\log|{\phi}(M(z;a_{j},b_{j}))|\geq\int_{0}^{y}\operatorname{Im}\mathfrak{g}^{\prime}(x+\mathrm{i}y){\mathop{}\!\mathrm{d}y}$ $\geq\left(c^{\prime}-\sqrt{y}\frac{2C^{\prime\prime}}{3{c^{\prime\prime}}}\right)\log|{\phi}(M(z;a_{j},b_{j}))|,$
from which the claim follows. Similar considerations work for $\operatorname{Im}y<0$, establishing the claim for $a_{j}\leq x\leq b_{j}$. Continuity of the ratio $\operatorname{Re}\mathfrak{g}(z)/\log|{\phi}(M(z;a_{j},b_{j}))|$ near $a_{j},b_{j}$ allows the desired bound to hold in neighborhoods of $a_{j},b_{j}$, completing the proof.
@@ Appendix D Hyperelliptic Riemann surfaces
Classic references for what follows are [9, 16, 5]. Much of what follows here was closely adapted from that in [18]. Using the intervals $a_{j},b_{j}],1\leq j\leq g+1,$ we define a Riemann surface by the zero locus of
@@
$\oint_{{\mathfrak{a}}_{i}}\sum_{k=1}^{g}c_{k}\mathop{}\!\mathrm{d}\nu_{k}=\sum_{k=1}^{g}c_{k}A_{ik}=\mathbf{e}_{i}^{T}\mathbf{A}\mathbf{c}=\mathbf{e}_{i}^{T}\mathbf{e}_{j}=\delta_{ij}.$
@@
$\frac{\theta\left(\pm\mathbf{u}(z)+{\boldsymbol{\tau}}\mathbf{e}_{j}+\mathbf{v}-\mathbf{d};{\boldsymbol{\tau}}\right)}{\theta\left(\pm\mathbf{u}(z)+{\boldsymbol{\tau}}\mathbf{e}_{j}-\mathbf{d};{\boldsymbol{\tau}}\right)}=\,\mathrm{e}^{\pm{v_{j}}}\frac{\theta\left(\pm\mathbf{u}(z)+\mathbf{v}-\mathbf{d};{\boldsymbol{\tau}}\right)}{\theta\left(\pm\mathbf{u}(z)-\mathbf{d};{\boldsymbol{\tau}}\right)},\quad{z\in[b_{j},a_{j+1}].}$
@@
where $\mathcal{C}$ is a clockwise-oriented simple contour that encircles $a_{1},b_{g+1}]$. Because all the differentials $\mathop{}\!\mathrm{d}\omega_{j}$ are of the form $P(z)/R(z)$ where $P$ is a degree g1g-1 polynomial and $R(z)=\mathrm{O}(z^{g+1})$ as $z\to\infty$, it follows that $\oint_{\mathcal{C}}\mathop{}\!\mathrm{d}\omega_{j}=0$. Thus, ignoring any poles $\boldsymbol{\Theta}$ may have, we find that $\boldsymbol{\Theta}$ satisfies the following jump conditions:
$\boldsymbol{\Theta}^{+}(z)=\begin{cases}\boldsymbol{\Theta}^{-}(z)\begin{bmatrix}0&1\\ 1&0\end{bmatrix}&z\in(a_{j},b_{j}),\vskip 3.0pt\\ \boldsymbol{\Theta}^{-}(z)\begin{bmatrix}\,\mathrm{e}^{-v_{j}}&0\\ 0&\,\mathrm{e}^{v_{j}}\end{bmatrix}&z\in(b_{j},a_{j+1}),\\ \boldsymbol{\Theta}^{-}(z)&z\in(-\infty,a_{1})\cup(b_{g+1},\infty).\end{cases}$
@@
From [21] (see also [36, Lemma 11.10]), these divisors are nonspecial so that the $\theta$ functions we consider do not vanish identically.
@@
It is important to note that (D.5) was first used in [14] and subsequently by many, see [13, 16, Chen2008a], for example.
@@ Appendix E Estimates of some exponential integrals
$E_{\rho}$ $=\{z~:~|z+\sqrt{z+1}\sqrt{z-1}|<\rho\},\quad\rho>1,$
@@
$I_{m}:=\int_{E_{\rho}}$ $|z^{2}-1|^{-\gamma}|{\phi(z)}|^{-m}\mathop{}\!\mathrm{d}A(z).$
We then set $z=1/2(w+w^{-1})$ for $1<|w|<\rho$ and change variables. We have
@@
$I_{m}=\int_{1<|w|<\rho}|w|^{2\gamma-4-m}2^{{2}\gamma-2}|w^{2}-1|^{2-2\gamma}\mathop{}\!\mathrm{d}A(w).$
@@
Now, suppose $0{\geq}2-2\gamma=-\sigma>-1$. We write $w=\rho\,\mathrm{e}^{\mathrm{i}\theta}$ and $\theta\in[-\pi/2,\pi/2]$, following [40], write
$|w-1|=\sqrt{(\rho-1)^{2}+4\rho\sin^{2}\left(\theta/2\right)}\geq C\left[(\rho-1)+|\theta|\right],$
for some $C>0$. Consider
$\int_{1}^{\rho}\int_{-\pi/2}^{\pi/2}\rho^{-m}|w-1|^{-\sigma}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\theta\leq C^{-\sigma}\int_{1}^{\rho}\rho^{-m}\left(\int_{-\pi/2}^{\pi/2}\frac{\mathop{}\!\mathrm{d}\theta}{((\rho-1)+|\theta|))^{\sigma}}\right)\mathop{}\!\mathrm{d}\rho$
For $\sigma<1$, the $\theta$ integral produces a continuous function of $\rho$. Since it therefore must be bounded, we can conclude the integral is $O(m^{-1})$. This argument can then be used for $w$ in the right and left half-planes separately, bounding $|w\pm 1|^{-\sigma}\leq 1$ for $\pm\mathrm{Re}\,w\geq 0$ to also conclude (E.1) for $\gamma<3/2$.